A castle in the faraway land of Aluossim was surrounded by four moats. One day, the castle was attacked and captured by a fierce tribe from the north. Guards were stationed at each bridge. Prince Juan was allowed to take a number of bags of gold from the castle as he went into exile. However, the guard at the first bridge demaded half of the bags of gold plus one more bag. Juan met his demand and proceeded to the next bridge. The guards at the second, third, and fourth bridges made identical demands, all of which the prince met. When Juan finally crossed all the bridges, a single bag of gold was left. With how many bags did Juan start? (hint:work backwards)
Answer by jorel1380(3719)

Working this one backwards, we get the pattern to be ((n+1)x2). We end with one single bag of gold. Before we had that we had((1+1)x2),or 4 bags of gold. Doing this again we get (4+1)x2 or 10 bags. (10+1)x2=22, (22+1)x2=46. So we started with 46 bags of gold.